OpenStudy (shubhamsrg):

If x, y, and k are positive numbers such that ((10x)/(x+y)) + ((20y)/(x+y)) = k and if x < y, which of the following could be the value of k? A. 10 B. 12 C. 15 D. 18 E. 30

OpenStudy (shubhamsrg):

nops..30 is not the ans..

OpenStudy (shubhamsrg):

what makes you go for 30?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

12 or 18

OpenStudy (anonymous):

is dat rite??

OpenStudy (shubhamsrg):

18's the ans.. but i'd to know your solns.. i'll post mine after some time..

OpenStudy (anonymous):

 in the LHS the second term is positive....so the ans is GREATER than 10 and ti get 15....(y/x+y)=1/2....which means x=y which is not possible and obviously 30 is ruled out...